Table 1. Return on Assets

Asset GBP

Budgeted profits for the year 10,575,000

Asset base at the start of the year 58,925,000

Budgeted asset base at the end of the year 69,500,000

Average asset base 64,212,500

ROA (using average of opening and closing assets) 16.5%


Calculate the budgeted ROA for sales outputs of 4,200,000 and 4,500,000 units. Assume that any additional profits over the original budget will be invested in the business and therefore increase the budgeted closing asset base.
1
Expert's answer
2020-05-20T10:13:01-0400

"ROA=\frac{net\ incom}{average\ total\ assets}"

x-profit for one unit or ROA for unit

y-quantity of sales

"profit=x*y"

"ROA=\frac{x*y}{average\ total\ assets}"


"16.5=\frac{x*4,200,000}{\frac{58,925,500+x*4,200,000}{2}}*100"

"0.165=\frac{2x*4,200,000}{58,925,000+x*4,200,000}"

"9,722,625+693,000x=2x*4,200,000"

"2.31=2x-0.165x"

"x=1.26"



"0.165*(58,925,000+x*4,500,000)=2x*4,500,000"

"9,722,625+742,500x=2x*4,500,000"

"2.16=2x-0.165x"

"x=1.18"